### v1

@@ 4.2 Discrimination Variables
To effectively discriminate neutron and $\gamma$ events from the AmBe source, different discrimination variables are considered as an alternative to the PSD technique. Fractions of integrated charges from different regions in time and cumulants ($C_{i}$) are calculated from the waveform for neutron and $\gamma$ as shown in Fig. 1(b). These variables are used as input for ML algorithms. We have considered the ratio of the integrated charge in the time window of 48 ns to 68 ns to the total charge in the time window of 46 ns to 250 ns. This ratio is denoted as head charge ($R_{head}$ [30]) variable and is given in Eq.(1).
@@
$k\sigma^{2}$ $=\frac{C_{4}}{C_{2}}$ (3)
The kurtosis by definition is a measure of relative tailedness of a distribution, which is a widely used parameter to study fluctuations in the hadron to quark gluon plasma phase transition by looking at the multiplicity distribution of produced particles in high energy heavy ion collisions [31, 32]. When multiplied by the variance, it enhances the magnitude of the fluctuations arising from the differences in the tail part of the distribution. As neutrons have a comparatively longer tail part, thus, $k\sigma^{2}$–given by Eq.(3)– is considered as a suitable input variable for the discrimination using ML algorithms.
Figure 4: Input variables used for the ML algorithms are (a) $R_{head}$, (b) $\langle t\rangle_{Q}$, and (c) $k\sigma^{2}$. The distributions are shown for neutron (red dashed) and $\gamma$ (blue) events.
Figure 4 (a), (b), and (c) shows the distribution of $R_{head}$, $\langle t\rangle_{Q}$, and $\kappa\sigma^{2}$ for neutron and $\gamma$ events, respectively. A reasonable degree of separation among different variables is observed for neutron and $\gamma$ events and hence can be used for training purposes of the ML algorithms.
@@ 4.3 Training of dataset
For the supervised ML algorithms, the training of models require an input dataset with labeled signal and background events. To construct training dataset, a data driven PSD based model is considered for initial labeling. Figure 5 (a) to (f) shows the projection of PSD distribution in few selected different integrated charge ranges. The mean ($\mu$) and the standard deviation ($\sigma$) for neutron and $\gamma$ events, as a function of charge, are obtained for each charge range by fitting the PSD projection with a double Gaussian function. Figure 6 (a)-(d) shows the variation of the $\mu_{n,\gamma}$ and the $\sigma_{n,\gamma}$ of neutron and $\gamma$ events as a function of charge from 200 to 4000 ADC. The variation is parameterized by fitting the data with a third-order polynomial function and the interpolation of this function is used to obtain $\mu$ and $\sigma$ at lower charge region where the neutron and $\gamma$ peaks cannot be fitted explicitly.
@@
Since the $\mu$ and $\sigma$ values show a variation as a function of integrated charge, we obtain a normalized variable, namely “$\mathrm{kSigma}$”, for the selection of neutron and $\gamma$ input events for training of ML algorithms. The “$\mathrm{kSigma}$” variable is calculated using Eq.(4), where the $k$ represent the number of standard deviation around the mean ($\mu$) of PSD distribution which is adapted dynamically across different charge regions for the selection of training events.
$\mathrm{(kSigma)_{n,\gamma}}$ $=\frac{PSD-\mu_{n,\gamma}}{\sigma_{n,\gamma}}$ (4)
@@
The training dataset contains neutron and $\gamma$ events selected on the basis of $\mathrm{(kSigma)_{n,\gamma}}$ parameter. Those events which satisfy the condition of $\mathrm{(kSigma)_{\gamma}}$ 3.0 are labeled as $\gamma$ events, where as events with $\mathrm{(kSigma)_{n}}$3.0 are labeled as neutron events. For each ML algorithm, a training dataset is prepared by randomly selecting 20,000 events each for neutron and $\gamma$ based on the above selection. Another dataset, containing 100,000 neutron and $\gamma$ events are used for the testing purpose of the ML algorithms.
@@
For MLPBNN, the first three hidden layers consist of N+4, N+4, and N+2 nodes where N represents the number of variables. Apart from the three input variable nodes, MLPBNN uses a bias node. A bias node provides a constant input to the nodes in the next layer and have their own weights associated with each connection which is also learned during training. The activation function in our case is a non linear sigmoid function which is represented by $\mathrm{tanh~{}(x)}$. This function introduces a nonlinearity into the network, allowing it to learn complex patterns in the data. In order to reduce the number of iterations to optimize the computation time an alternative approach called the Broyden-Fletcher-Goldfarb-Shannon (BFGS) method is utilized while adapting the synapse weight. The combination of this particular architecture is found to be optimal in terms of performance and convergence of the network. The MLPBNN is trained for 200 epochs using the BFGS optimization method for the convergence of the algorithm. In parallel, the SVM algorithm is tested with Radial Basis Function (RBF) kernel mode. The RBF kernel is usually used when the data are non linear in nature. Apart from the calculation of the squared distances between points, it has the $\Gamma$ parameter which controls how well model fits the data. The RBF mode used in our training dataset has $\Gamma$ parameter value of 0.25 and with a tolerance of 0.001.
@@ 4.4 Response from MLPBNN and SVM
Figure  7 (a) shows the comparison of the Receiver Operating Characteristic (ROC) curve for the MLPBNN and SVM algorithm. The background rejection efficiency or False Positive Rate (FPR) is represented as a function of the signal selection efficiency or the True Positive Rate (TPR). From the ROC curve, MLPBNN is found to achieve better background rejection for the same signal efficiency as compared to SVM. Also shown, Fig. 7 (b) and (c), are the classifier responses for SVM and MLPBNN methods, respectively. A reasonable level of discrimination between the neutron and $\gamma$ events is achieved from the ML algorithms with our variables and optimized parameters.
@@ 5 Time of Flight Experiment
The trigger scheme used for the ToF measurement in the waveform digitizer is configured such that a start signal timestamp is generated when the $\mathrm{CeBr_{3}}$ detector registers a 4.4 MeV $\gamma$. The stop signal timestamp is recorded when either a neutron or a $\gamma$ event is triggered in the PS detector. To reduce the rate of accidental background events, a correlated time window of 200 ns interval between start and stop time is used for the data acquisition. Due to the relatively small active area of the PS detector, the overall trigger acceptance of the setup is relatively low and the data presented for the ToF analysis consists of 3 days. The raw waveforms from both detectors are recorded and the PS detector data is analyzed in the same manner as for the singles events. The ToF variable is calculated by taking the the recorded start and stop timestamp generated from $\mathrm{CeBr_{3}}$ and PS detector, respectively. A clear separation between the 650 extracted neutron and $\gamma$ events is observed in the ToF distribution, as shown in Fig. 10(b). The events with $\mathrm{ToF>30}$ ns are tagged as neutron events (red hatched region), while events with $\mathrm{ToF<25}$ns are tagged as $\gamma$ events (blue hatched region). The PSD parameter is then calculated for these tagged events, in the same manner as done for singles events, and are compared with the PSD distribution obtained from the discrimination of neutron and $\gamma$ events using the ML score.
@@ 6 Discussion
Figure 13: Confusion Matrix for (a) SVM and (b) MLPBNN.
To quantify the robustness of our ML algorithms, we calculate the confusion matrix as shown in Fig. 13 (a) and (b), which represents the discrimination performance of the SVM and MLPBNN models, respectively. The true positives (TP) define the actual $\gamma$ events correctly classified as $\gamma$, and the false positives (FP) correspond to the neutron events incorrectly predicted as $\gamma$ events. Similarly, the true negatives (TN) defines the actual neutron events which are correctly classified as neutron events and false negative (FN) corresponds to the $\gamma$ events incorrectly classified as neutron events. The true neutron and $\gamma$ events are obtained from ToF results and the predicted events are from the ML algorithms. From this formulation, we compute some of the key measures for the MLPBNN and SVM algorithms which are given in the Table. 2. The accuracy, defined as the ratio of sum of TP and TN to the total number of events used for the classification. It gives the overall performance of the algorithm in discrimination of the neutron and $\gamma$ events. Precision is defined as the ratio of TP and the sum of TP and FP, and represents the quality of the models positive predictions. Similarly recall is defined as the ratio of TP and the sum of TP and FN, measuring the probability of predicting positives. Specificity is another important parameter which evaluates the classification of the models particularly in binary classification as in our case. It measures the ability of a model to correctly identify the negative instances and is calculated by taking the ratio of TN to the sum of TN and FP.
@@
MLPBNN 0.80 0.79 0.84 0.77 0.81 0.24 0.16
SVM 0.79 0.80 0.81 0.76 0.80 0.24 0.19
Lastly, the F1-score provides a measure of performance of the model particularly when there is an imbalance of events in the datasets. The F1-score is represented by $\mathrm{2\times Precision\times Recall/(Precision+Recall)}$. Apart from these performance metric to evaluate our ML algorithms, we also define the type-I and type-II errors. Type-I error is calculated as ratio of FP to the sum of FP and TN and effects the precision of a model which measured accuracy of positive predictions. Similarly, type-II errors calculated from the ratio of FN to the sum of TP and FN denotes the impact on the recall of a model by signifying how well the model can identify all the actual predictions. Overall, the key metric for the ML algorithms used for the neutron and $\gamma$ discrimination are reasonable and shows promising future for the use of such algorithms for the particle identification in lower energy regions.
@@ 7 Conclusion
In this study, machine learning based algorithms are used on signals obtained from plastic scintillator detector with PSD capabilities. To enhance the discrimination of neutron and $\gamma$ events in the low energy regime where the traditional PSD technique has inferior performance, MLPBNN and SVM, are used to separate them using raw waveforms. The variables, $\mathrm{R_{head}}$, $\langle t\rangle_{Q}$, and $\mathrm{\kappa\sigma^{2}}$, calculated directly from the waveform, show promising improvements in the discrimination of neutron and $\gamma$ events at lower energies beyond 151 $keV_{ee}$. Additionally, to cross-check the validity of the discrimination obtained from machine learning algorithms, a comparison of results for neutron and $\gamma$ events is done with the results obtained from a time-of-flight experiment. The time-of-flight experiment is particularly effective in the low energy regime, hence providing a platform to benchmark the ML algorithms. The results from ML score and from ToF data shows an reasonable level of separation of neutron and $\gamma$ events in the low energy regime. Results obtained from the ML algorithms are consistent with those obtained from the ToF experiment, with an accuracy of 0.80 and 0.79 for MLPBNN and SVM, respectively. From the ROC curve as well as the sensitivity test of ML models, MLPBNN is found to have better discriminatory power in all charge regions. For future studies, we aim to explore more advanced ML techniques such as Convolution Neural Network (CNN) and Deep Neural Network (DNN) to access the discrimination in even lower energy ranges with the improvements in lowering of the detector and electronic thresholds from the plastic scintillator detectors.

### v2

@@ 4.2 Discrimination Variables
To effectively discriminate neutron and $\gamma$ events from the AmBe source, different discrimination variables are considered as an alternative to the PSD technique. Fractions of integrated charges from different regions in time and cumulants ($C_{i}$) are calculated from the waveform for neutron and $\gamma$ as shown in Fig. 1(b). These variables are used as input for ML algorithms. We have considered the ratio of the integrated charge in the time window of 48 ns to 68 ns to the total charge in the time window of 46 ns to 250 ns. This ratio is denoted as head charge ($R_{head}$ [Lee:2024lvr]) variable and is given in Eq.(1).
@@
$\kappa\sigma^{2}$ $=\frac{C_{4}}{C_{2}}$ (3)
The kurtosis by definition is a measure of relative tailedness of a distribution, which is a widely used parameter to study fluctuations in the hadron to quark gluon plasma phase transition by looking at the multiplicity distribution of produced particles in high energy heavy ion collisions [STAR:2010mib, Garg:2013ata]. When multiplied by the variance, it enhances the magnitude of the fluctuations arising from the differences in the tail part of the distribution. As neutrons have a comparatively longer tail part, thus, $\kappa\sigma^{2}$–given by Eq.(3)– is considered as a suitable input variable for the discrimination using ML algorithms.
Figure 4: Input variables used for the ML algorithms are (a) $R_{head}$, (b) $\langle t\rangle_{Q}$ (in ns), and (c) $\kappa\sigma^{2}$. The distributions are shown for neutron (red dashed) and $\gamma$ (blue) events.
Figure 4 (a), (b), and (c) shows the distribution of $R_{head}$, $\langle t\rangle_{Q}$(in ns), and $\kappa\sigma^{2}$ for neutron and $\gamma$ events, respectively. A reasonable degree of separation among different variables is observed for neutron and $\gamma$ events and hence can be used for training purposes of the ML algorithms.
@@ 4.3 Training of dataset
For the supervised ML algorithms, the training of models require an input dataset with labeled signal and background events. To construct training dataset, a data driven PSD based model is considered for initial labeling. Figure 5 (a) to (f) shows the projection of PSD distribution in few selected different integrated charge ranges. The mean ($\mu$) and the standard deviation ($\sigma$) for neutron and $\gamma$ events, as a function of charge, are obtained for each charge range by fitting the PSD projection with a double Gaussian function. Figure 6 (a)-(d) shows the variation of the $\mu_{n,\gamma}$ and the $\sigma_{n,\gamma}$ of neutron and $\gamma$ events as a function of integrated charge from 200 to 4000. The variation is parameterized by fitting the data with a third-order polynomial function and the interpolation of this function is used to obtain $\mu$ and $\sigma$ at lower charge region where the neutron and $\gamma$ peaks cannot be fitted explicitly.
@@
Since the $\mu$ and $\sigma$ values show a variation as a function of integrated charge, we obtain a normalized variable, namely $\mathrm{kSigma}$”, for the selection of neutron and $\gamma$ input events for training of ML algorithms. The $\mathrm{kSigma}$” variable is calculated using Eq.(4), where the $k$ represent the number of standard deviation around the mean ($\mu$) of PSD distribution which is adapted dynamically across different charge regions for the selection of training events.
$(\mathrm{kSigma})_{n}$ $=\frac{PSD-\mu_{n}}{\sigma_{n}},\qquad(\mathrm{kSigma})_{\gamma}=\frac{PSD-\mu_{\gamma}}{\sigma_{\gamma}}.$ (4)
Figure 7: $\mathrm{(kSigma)_{n}}$ distribution is shown in different integrated charge bins. The combined fit (in teal) as well as $\gamma$ and neutron fit is shown in blue and red Gaussians, respectively. The shaded blue region represents the $\gamma$-events labeled for training.
Figure 8: $\mathrm{(kSigma)_{\gamma}}$ distribution is shown in different integrated charge bins. The combined fit (in teal) as well as $\gamma$ and neutron fit is shown in blue and red Gaussians, respectively. The shaded red region represents the neutron-events labeled for training.
The training dataset contains neutron and $\gamma$ events selected on the basis of $\mathrm{(kSigma)_{n,\gamma}}$ parameter. Figure 7 shows the $\mathrm{(kSigma)_{n}}$ distribution in different integrated charge bins. Also shown, are the combined Gaussian fit of the distributions along with the individual components representing the neutron and $\gamma$ event fractions. The events satisfying condition $\mathrm{(kSigma)_{n}}$ ¡ -3.0, shown in blue shaded region, represents the selected true label for training of $\gamma$ events. Similarly, Fig. 8 shows the $\mathrm{(kSigma)_{\gamma}}$ distribution in various integrated charge bins along with the Gaussian fits. The red shaded region in the Fig. 8 (a)-(f) shows the selection of neutron events for the true label with the selection criteria of $\mathrm{(kSigma)_{\gamma}}$ ¿ 3.0. With this selection criteria on $\mathrm{(kSigma)_{n}}$ and $\mathrm{(kSigma)_{\gamma}}$ for $\gamma$ and neutron true labels for training of ML algorithms, the uncertainty in the selection of true labels is reduced to less than $0.1\%$ in the studied integrated charge range.
@@
For each ML algorithm, a training dataset is prepared by randomly selecting 20,000 events each for neutron and $\gamma$ based on the above selection. Another dataset, containing 100,000 neutron and $\gamma$ events are used for the testing purpose of the ML algorithms. To identify the optimal configuration, for both MLPBNN and SVM algorithms, optimization is done by tuning the set of input parameter dependencies and stability of the algorithms on the variables of interest.
@@
For MLPBNN, the convergence of the algorithm is tested with different numbers of hidden layers.The first three hidden layers consist of N+4, N+4, and N+2 nodes where N represents the number of variables. Apart from the three input variable nodes, MLPBNN uses a bias node. A bias node provides a constant input to the nodes in the next layer and have their own weights associated with each connection which is also learned during training. The activation function in our case is a non linear sigmoid function which is represented by $\mathrm{tanh~(x)}$. This function introduces a nonlinearity into the network, allowing it to learn complex patterns in the data. In order to reduce the number of iterations to optimize the computation time an alternative approach called the Broyden-Fletcher-Goldfarb-Shannon (BFGS) method is utilized while adapting the synapse weight. The combination of this particular architecture is found to be optimal in terms of performance and convergence of the network. The MLPBNN is trained for 200 epochs using the BFGS optimization method. The number of epochs are optimized by monitoring the performance of the algorithm and by comparing the outputs from training and testing datasets to check underfitting or overfitting of the results. Additionally, the degree of correlation between the variables used for training was optimized to minimize the potential bias. The MLPBNN provides a method for estimating the relative importance of each input variable through a variable ranking technique. This method utilizes the sum of the squared weights of the connections between the input neuron corresponding to variable $i$ and all neurons in the first hidden layer. The importance $I_{i}$ of the input variable $i$ is defined as,
$I_{i}=x_{i}^{2}\sum_{j=1}^{n_{h}}\left(w_{ij}^{(1)}\right)^{2},\quad i=1,\dots,n_{\text{var}}$ (5)
where, $x_{i}$ is the sample mean of the input variable $i$, $w_{ij}^{(1)}$ is the weight connecting input neuron $i$ to hidden neuron $j$ in the first hidden layer, $n_{h}$ is the number of neurons in the first hidden layer, and $n_{\text{var}}$ is the total number of input variables. The importance of the three input variables are presented in the table 1. It can be seen that the input parameter $\kappa\sigma^{2}$ provides better ranking in terms of separability than other variables.
Table 1: Ranking of input variables for MLPBNN.
Rank Variable Importance (in%)
1 $\kappa\sigma^{2}$ 61.03
2 $\langle t\rangle_{Q}$ 47.28
3 $R_{head}$ 43.76
In parallel, the SVM algorithm is tested with Radial Basis Function (RBF) kernel mode. The RBF kernel is usually used when the data are non linear in nature. Apart from the calculation of the squared distances between points, it has the $\Gamma$ parameter which controls how well model fits the data. For SVM algorithm, the tolerance and $\Gamma$ values are tuned to minimize the false positive rate which is given by Type-I error. The RBF mode used in our training dataset has a $\Gamma$ value of 0.25 and with a tolerance of 0.001. Though higher $\Gamma$ (=1) value yields marginal gain ($0.25\%$) in area under the curve, but it introduces an overfitting risk. Thus we have chosen a value for which the decision boundary is smooth enough to classify efficiently while preserving the model from over-fitting.
@@ 4.4 Response from MLPBNN and SVM
Figure  9 (a) shows the comparison of the Receiver Operating Characteristic (ROC) curve for the MLPBNN and SVM algorithm. The background rejection efficiency or False Positive Rate (FPR) is represented as a function of the signal selection efficiency or the True Positive Rate (TPR). From the ROC curve, MLPBNN is found to achieve better background rejection for the same signal efficiency as compared to SVM. Also shown, Fig. 9 (b) and (c), are the classifier responses for SVM and MLPBNN methods, respectively. A reasonable level of discrimination between the neutron and $\gamma$ events is achieved from the ML algorithms with our variables and optimized parameters.
@@ 5 Time of Flight Experiment
$\mathrm{E_{n}=\frac{1}{2}m_{n}v^{2}=\frac{1}{2}m_{n}\left(\frac{L}{ToF}\right)^{2}},$ (6)
where, $\mathrm{m_{n}}$ is the mass of the neutron, $\mathrm{v}$ is the velocity of the neutron, $\mathrm{L}$ is the flight path length, and $\mathrm{ToF}$ is the time-of-flight. For 1 meter flight path, time-of-flight for neutron and $\gamma$ events is given by,
$\text{ToF}_{n}=\frac{72.3}{\sqrt{E_{n}}},\quad\text{ToF}_{\gamma}=\frac{1}{c}.$ (7)
@@
Figure 13: Neutron energy distribution, calculated from ToF, for different ToF bins having an integrated charge in 200 to 800.
The trigger scheme used for the ToF measurement in the waveform digitizer was configured such that a start signal timestamp was generated when the $\mathrm{CeBr_{3}}$ detector registers a 4.4 MeV $\gamma$ event. The stop signal timestamp was recorded when either a neutron or a $\gamma$ event was triggered in the PS detector. To reduce the rate of accidental background events, a correlated time window of 200 ns ($\mathrm{E_{n}\sim 130}$ keV) interval between start and stop time was used for the data acquisition. Due to the relatively small active area of the PS detector, the overall trigger acceptance of the setup was relatively low and the data presented for the ToF analysis consists of 3 days. The raw waveforms from both the detectors were recorded and the PS detector data was analyzed in the same manner as for the singles events. The ToF variable was calculated by taking the recorded start and stop timestamp generated from $\mathrm{CeBr_{3}}$ and PS detector, respectively. In the charge range 200 to 4000, the ToF distribution shows a clear separation between the extracted neutron and $\gamma$ events, as shown in Fig. 12(b). The events with $\mathrm{ToF>30}$ ns are tagged as neutron events (red hatched region), while events with $\mathrm{ToF<25}$ ns are tagged as $\gamma$ events (blue hatched region). For the integrated charge range 200 to 800, a total of 650 $\gamma$ and neutron events are filtered. For the same charge range, Fig. 13 shows the neutron energy distribution, calculated from Eq. 7. The energy deposition by a recoiling proton is affected by the finite detector acceptances, charge collection inefficiencies, and quenching effects arising from non-linear energy deposition. Consequently, a fraction of neutron events ($\sim 36\%$) are found to have proton recoil energy deposition in the lower $\mathrm{E_{dep}}$ range of 150 to 400 $\mathrm{keV_{ee}}$ where the conventional PSD technique fails to discriminate neutrons and $\gamma$-rays. The PSD parameter is then calculated for these tagged events, in the same manner as done for singles events, and are compared with the PSD distribution obtained from the discrimination of neutron and $\gamma$ events using the ML score.
@@ 6 Discussion
Figure 16: Confusion Matrix for (a) SVM, (b) MLPBNN, and (c) CCM is shown for the integrated charge range 200-800. The yellow quadrants represent the correctly classified events (True Positives and True Negatives), and the blue quadrants report the misclassified events (False positives and False Negatives).
To quantify the robustness of our ML algorithms, we calculate the confusion matrix as shown in Fig. 16 (a), (b), and (c), which represents the discrimination performance of the SVM, MLPBNN, and CCM, respectively. The true labels in the x-axis are obtained from the ToF technique and the y-axis represents the predicted labels from ML algorithms or PSD. The PSD-based selection of neutron and $\gamma$ events in the integrated charge range 200-800, is motivated from the well separated bands in the higher charge range. The true positives (TP) define the actual $\gamma$ events correctly classified as $\gamma$, and the false positives (FP) correspond to the neutron events incorrectly predicted as $\gamma$ events. Similarly, the true negatives (TN) defines the actual neutron events which are correctly classified as neutron events and false negative (FN) corresponds to the $\gamma$ events incorrectly classified as neutron events. From this formulation, we compute some of the key measures for the MLPBNN, SVM algorithms, and CCM which are given in the Table. 3. The accuracy, defined as the ratio of sum of TP and TN to the total number of events used for the classification. It gives the overall performance of the algorithm in discrimination of the neutron and $\gamma$ events. Precision is defined as the ratio of TP and the sum of TP and FP, and represents the quality of the models positive predictions. Similarly recall is defined as the ratio of TP and the sum of TP and FN, measuring the probability of predicting positives. Specificity is another important parameter which evaluates the classification of the models particularly in binary classification as in our case. It measures the ability of a model to correctly identify the negative instances and is calculated by taking the ratio of TN to the sum of TN and FP.
@@
MLPBNN 0.80 0.84 0.79 0.81 0.81 0.19 0.21
SVM 0.79 0.81 0.80 0.78 0.80 0.22 0.20
CCM 0.46 0.88 0.02 0.99 0.04 0.01 0.98
Lastly, the F1-score provides a measure of performance of the model particularly when there is an imbalance of events in the datasets. The F1-score is represented by $\mathrm{2\times Precision\times Recall/(Precision+Recall)}$. Apart from these performance metric to evaluate our ML algorithms, we also define the type-I and type-II errors. Type-I error is calculated as ratio of FP to the sum of FP and TN and effects the precision of a model which measured accuracy of positive predictions. Similarly, type-II errors calculated from the ratio of FN to the sum of TP and FN denotes the impact on the recall of a model by signifying how well the model can identify all the actual predictions. In Fig. 16, the yellow entries indicate the correctly classified entries, contributing to an overall accuracy of $80\%$ and $79\%$ for MLPBNN and SVM, respectively. Similarly, the blue entries representing falsely classified entries FP and FN, contribute to recall and precision, respectively. Overall, the key metrics for the ML algorithms used for the neutron and $\gamma$ discrimination are reasonable and shows better discrimination performance as compared to the CCM. This hints a promising future for the use of such algorithms for the particle identification in lower energy deposition regions.
@@ 7 Conclusion
In this study, machine learning based algorithms are used on signals obtained from plastic scintillator detector with PSD capabilities. To enhance the discrimination of neutron and $\gamma$ events in the low energy regime where the traditional PSD technique has inferior performance, MLPBNN and SVM, are used to separate them using raw waveforms. The variables, $\mathrm{R_{head}}$, $\langle t\rangle_{Q}$, and $\mathrm{\kappa\sigma^{2}}$, calculated directly from the waveform, show promising improvements in the discrimination of neutron and $\gamma$ events at lower energies beyond 151 $\mathrm{keV_{ee}}$. Additionally, to cross-check the validity of the discrimination obtained from machine learning algorithms, a comparison of results for neutron and $\gamma$ events is done with the results obtained from a time-of-flight experiment. The time-of-flight experiment is particularly effective in the low energy regime, hence providing a platform to benchmark the ML algorithms. The results from ML score and from ToF data shows an reasonable level of separation of neutron and $\gamma$ events in the low energy regime. Results obtained from the ML algorithms are consistent with those obtained from the ToF experiment, with an accuracy of 0.80 and 0.79 for MLPBNN and SVM, respectively. From the ROC curve as well as the sensitivity test of ML models, MLPBNN is found to have better discriminatory power in all charge regions. For future studies, we aim to explore more advanced ML techniques such as Convolution Neural Network (CNN) and Deep Neural Network (DNN) to access the discrimination in even lower energy ranges with the improvements in lowering of the detector and electronic thresholds from the plastic scintillator detectors.